OpenStudy (anonymous):

Solve 8sin2x – 6 = 0

OpenStudy (anonymous):

Oops its suppose to be \[8\sin^2x-6=0\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf 8sin2x – 6 = 0 }\) \(\large\color{black}{ \bf 8sin2x = 6 }\) \(\large\color{black}{ \bf sin2x = 6/8 }\) \(\large\color{black}{ \bf sin2x = 3/4 }\) \(\large\color{black}{ \bf sin(x+x) = 3/4 }\) \(\large\color{black}{ \bf sin(x)cos(x)+sin(x)cos(x) = 3/4 }\) \(\large\color{black}{ \bf 2sin(x)cos(x) = 3/4 }\) \(\large\color{black}{ \bf sin(x)cos(x) = 3/8 }\) \(\large\color{black}{ \bf sin^2(x)cos^2(x) = 9/64 }\) \(\large\color{black}{ \bf sin^2(x)(1-sin^2x) = 9/64 }\) \(\large\color{black}{ \bf sin^2(x)-sin^4x = 9/64 }\)

OpenStudy (solomonzelman):

Let sin^2x= a and solve for x.

OpenStudy (anonymous):

I am so confused!

OpenStudy (anonymous):

Which part do you get stuck at?

OpenStudy (anonymous):

After doing solomon's work you need to solve for x from the trinomial